A manufacturer buys motorbikes at a cost price of Rs 30600.He sells it to B,B sells it to C and C sells it to D.The rate of VAT is 12.5% find_( 1) total amount of VAT paid and, (2) the cost price for the buyer if profit at each stage of selling chain is Rs. 1000.

Grade:NULL

1 Answers

Arun
6 years ago
Dear Aayush
 

Selling price of motorbike to the manufacture = Rs 30600 + Rs 1,000 = Rs 31600

Amount of tax collected by manufacture A 

= 12.5% of Rs 31600

 = 12.5*31600/100 = 3950

 

Dealer B earns a profit of Rs 1000, so the value added by dealer B is Rs 1000.

Amount of VAT to be paid by B = 12.5% of Rs 1000 = 12.5*1000/100 = 125

 

Dealer C earns a profit of Rs 1000, so the value added by dealer C is Rs 1000.

Amount of VAT to be paid by C = 12.5% of Rs 1000 = 12.5 *1000/100 = 125

 

D earns a profit of Rs 1000, so the value added by D is Rs 1000.

Amount of VAT to be paid by D = 12.5% of Rs 1000 = 12.5*1000/100 = 125

 

Total amount of VAT paid = Rs 3950 + Rs 125 + Rs 125 + Rs 125 = Rs 4325

(2)

Value of motorbike paid by the consumer

= Rs 31600 + Rs 1000 + Rs 1000 + Rs 1000

= Rs 34600

Tax paid by the consumer 

= 12.5% of Rs 34,600

12.5 *34600/100 = 4325

∴ Amount paid by the consumer = Rs 34600 + Rs 4325 = Rs 38925

 

Regards

Arun (askIITians forum expert)